12/27/99 Day shift summary

Valeri Jejer (jejer@uvahea.phys.virginia.edu)
Mon, 27 Dec 1999 17:02:29 -0500 (EST)

Shift Summary for day shift, 27/Dec/99

On Shift: Ricardo, Valeri.

Hours of beam available at all intensities: 8.4
Hours of beam available at nominal intensity: 8.4
Hours of E799 data (special runs not included): 7.6
Hours of special runs (laser scans, muon runs...): 0.15

Accounting of lost time:

run starts/stops, fresh start, magnet polarity change.

========================================================================
Runs ended this shift:
Run 15304 - 14.0 M evts - BEAM_799
Run 15305 - 6.3 M evts - BEAM_799
Run 15306 - 0.35 M evts - BEAM_MU_4E12
Run 15307 - 10.6 M evts - BEAM_799

Tapes brought to FCC this shift:

(including tapes for online split)

UP(ABC)265-267, UPS089
========================================================================
Accesses during shift:

Under preparation as of the end of the shift.

========================================================================
Overview of shift:

Good beam all shift with intensity around 9e12 with only a couple of
empty spills and DA working well (only three kfrends caused by one
problem) except L3 monitoring problems caused by CsI channel 947 (ped.
exp >0).
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15304 7:12 L3 accept fraction alarms, filter planes disappear
from monitoring histograms one by one.

7:22 Stop the run and do fresh start.

15305 7:35 Start new run after fresh start.

8:00 Noticed that mass monitoring histograms are screwed
up, especially pi+pi-pi0Dal. Suspect CsI 947 which
has ped. exp > 0 and is shoing up on kdisp as
always having energy > 10 GeV. Etot monitoring is
also affected. L3 accept fraction is OK. By BobT's
request checked if L3 filter has mass cuts for any
modes and didn't find any.

10:25 Tapes are full, run stopped. Call MCR and ask them
to
switch us into muon mode. Flip NM4AN from NEG to
POS.

15306 10:30 Init and start muon run.

10:42 Stop muon run, ramp up the magnet and switch back to
799 mode.

15307 10:50 Start new run with old problems.

12:30 DYC53 is busy followed by two DC DYC busies, took 3
kfrends to fix.

15:28 Stop run because a bunch of people wants to go on
access to fix a bunch of problems.

========================================================================
Follow up Notes on Previous Shift
> Swap DPMT 947 during >1 hour down-time.
in progress.

========================================================================
Items to Follow up on Next Shift